# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### v2

@@ 2. $(2)$
where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ is defined as in (4.4) and (4.5).
@@
$(1)$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)- (4.2c) holds. We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since, $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1, $u\neq 0,$ and hence $\bar{u}\neq 0,$ which implies $\lambda$ is real, thus a contradiction to $\Im(\lambda)\neq 0.$
Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular. Again, since $Cv=0,$ from (4.2c), and similar to (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions
$\lambda^{\pm i}=\frac{2sq+p\pm i\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.26)
@@
where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds
@@
Again, considering the condition $q>(p^{2}/4),$ (4.27) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound
@@
On the other side, nonsingularity of $(\mu\Lambda_{1}-A)$ and (4.2a) enables us the following identity
@@
Consider the following eigenvalue decomposition of $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$
@@
where $\mathcal{D}=\mathrm{diag}(\theta_{i})\in{\mathbb{R}}^{n\times n}$ and $\theta_{i}>0,\,i=1,2,\ldots,n$ are in $\sigma(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})$ and $V\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix. Then, (4.33) can be written as
@@
$\Theta_{1}={\displaystyle{\mathrm{diag}\left(\frac{\Re(\mu)-\theta_{i}}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right),\Theta_{2}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right)}},$
from (4.34), and the fact $\Im(\mu)\neq 0,$ we get
@@
Hence, (4.39) yields the following bound
@@
$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}$ (4.43)
Combining (4.41) and (4.43) leads to the following bounds
@@
${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}(\frac{\xi_{\max}}{2}+\frac{1}{s})}$
@@
Let ${\mathscr{P}}_{SS}$ be defined as in (1.10) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then,
@@
Since ${\mathscr{P}}_{SS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then desired bounds will be obtained by setting $\Lambda_{1}=\Lambda_{2}=\Lambda_{3}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ into Theorem 4.3. $\blacksquare$
@@
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11), and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0,$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then $(1)$ If $Cv=0,$ $\lambda$ satisfies
@@
Since ${\mathscr{P}}_{EGSS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, then the desired bounds will be obtained by setting $\Lambda_{1}=\frac{1}{2}\alpha P,$ $\Lambda_{2}=\frac{1}{2}\beta Q$, $\Lambda_{3}=\frac{1}{2}\gamma W$ and $s=\frac{1}{2}$ in Theorem 4.3. $\blacksquare$
@@ 5 Local PESS (LPESS) preconditioner
$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-C^{T}\\ 0&C&\Lambda_{3}\end{matrix}\right],$
@@
where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $\frac{1}{s}$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
@@
Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ eigenvalue equal to $\frac{1}{s}.$ The remaining $m+p$ eigenvalues satisfies the following:
@@ 1. $(1)$
    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\theta_{\min}}{\vartheta_{\max}+s\theta_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
@@ 2. $(2)$
    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\theta_{\max}}{1+s\vartheta_{\min}+s^{2}\theta_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\theta_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ and $\theta_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-1}C^{T}).$
@@
where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&C^{T}\\ -C&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [28]. Therefore, by applying the Corollary 4.1 and Theorem 4.3 of [28], we obtain the desired bounds in $(1)$ and $(2).$blacksquare

### v3

@@ 2. $(2)$
where $\xi_{\max},\xi_{\min},\eta_{\max},\eta_{\min}$ and $\theta_{\max}$ are defined as in (4.4) and (4.5).
@@
$(1)$   Let $\lambda$ be an eigenvalue of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and the corresponding eigenvector is ${\bf p}=[u^{T},v^{T},w^{T}]^{T}.$ Then, the system of linear equations in (4.2a)-(4.2c) holds. We assert that, $v\neq 0,$ otherwise from (4.2a), we get $Au=\lambda(\Lambda_{1}+sA)u,$ which further implies $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}\bar{u}=\lambda\bar{u},$ where $\bar{u}=(\Lambda_{1}+sA)^{\frac{1}{2}}u.$ Since $(\Lambda_{1}+sA)^{-\frac{1}{2}}A(\Lambda_{1}+sA)^{-\frac{1}{2}}$ is a SPD matrix and by Lemma 4.1 $u\neq 0$ and thus we have $\bar{u}\neq 0.$ This implies $\lambda$ is real, leading to a contradiction to $\Im(\lambda)\neq 0.$
Initially, we consider the case $Cv=0.$ The aforementioned discussion suggests that $Au\neq\lambda(\Lambda_{1}+sA)u$ for any $u\neq 0.$ Consequently, $(\mu\Lambda_{1}-A)$ is nonsingular. Again, since $Cv=0,$ from (4.2c) and following a similar process as in (4.10)-(4.12), we obtain quadratic equation (4.13) in $\lambda$, which has complex solutions:
$\lambda^{\pm i}=\frac{2sq+p\pm i\sqrt{4q-p^{2}}}{2(s^{2}q+sp+1)}$ (4.26)
@@
where $p,q,s>0.$ Then $\Psi_{1}$ is monotonically increasing in the argument $q>0$ while decreasing in the argument $p>0.$ Due to the fact that $q\leq\lambda_{\max}(\Lambda_{1}^{-1}B^{T}\Lambda_{2}^{-1}B)=\eta_{\max},$ we get the following bounds:
@@
Again, considering the condition $q>(p^{2}/4),$ (4.27) adheres to the inequality $(p/(2+sp))\leq|\lambda^{\pm i}|.$ Notably, the function $\Psi_{2}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$ defined by $\Psi_{2}(p)=p/(2+sp),$ where $p,s>0,$ is monotonically increasing in the argument $p>0,$ we obtain the following bound:
@@
On the other side, nonsingularity of $(\mu\Lambda_{1}-A)$ and (4.2a) enables us the following identity:
@@
Consider the following eigenvalue decomposition of $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$:
@@
where $\mathcal{D}=\mathrm{diag}(\theta_{i})\in{\mathbb{R}}^{n\times n}$ and $\theta_{i}>0,\,i=1,2,\ldots,n,$ are in $\sigma(\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}})$ and $V\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix. Then, (4.33) can be written as
@@
$\Theta_{1}={\displaystyle{\mathrm{diag}\left(\frac{\Re(\mu)-\theta_{i}}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right),\leavevmode\nobreak\ \Theta_{2}=\mathrm{diag}\left(\frac{1}{(\Re(\mu)-\theta_{i})^{2}+\Im(\mu)^{2}}\right)}},$
from (4.34) and the fact $\Im(\mu)\neq 0,$ we get
@@
Hence, (4.39) yields the following bound:
@@
$\mathcal{X}\geq\frac{I}{\xi^{2}_{\max}+\eta_{\max}+\theta_{\max}}.$ (4.43)
Combining (4.41) and (4.43) leads to the following bounds:
@@
${\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\min}\eta_{\min}}{2(\xi_{\max}^{2}+\eta_{\max}+\theta_{\max})}+\frac{1}{s}\right)\leq\frac{1}{s}-\Re(\lambda)\leq\frac{1}{|\tau|^{2}}\left(\frac{\xi_{\max}}{2}+\frac{1}{s}\right)}$
@@
Let ${\mathscr{P}}_{SS}$ be defined as in (1.12) with $\alpha>0$ and let $\lambda$ be a non-real eigenvalue of the SS preconditioned matrix ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then,
@@
Since ${\mathscr{P}}_{SS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1, then desired bounds will be obtained by setting $\Lambda_{1}=\Lambda_{2}=\Lambda_{3}=\frac{1}{2}\alpha I$ and $s=1/2$ into Theorem 4.3. $\blacksquare$
@@
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.13) and let $P,Q$ and $W$ are SPD matrices with $\alpha,\beta,\gamma>0.$ Let $\lambda$ be a non-real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0$ and ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the corresponding eigenvector. Then $(1)$ If $Cv=0,$ $\lambda$ satisfies
@@
Since ${\mathscr{P}}_{EGSS}$ is a special case of ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1, then the desired bounds will be obtained by setting $\Lambda_{1}=\frac{1}{2}\alpha P,$ $\Lambda_{2}=\frac{1}{2}\beta Q$, $\Lambda_{3}=\frac{1}{2}\gamma W$ and $s=1/2$ in Theorem 4.3. $\blacksquare$
@@ 5 Local PESS (LPESS) preconditioner
$\mathfrak{L}=\left[\begin{matrix}I&0&0\\ -BA^{-1}&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{U}=\left[\begin{matrix}I&A^{-1}B^{T}&0\\ 0&I&0\\ 0&0&I\end{matrix}\right],\mathfrak{D}=\left[\begin{matrix}A&0&0\\ 0&Q&-C^{T}\\ 0&C&0\end{matrix}\right],\widetilde{\mathfrak{D}}=\left[\begin{matrix}sA&0&0\\ 0&\Lambda_{2}+sQ&-sC^{T}\\ 0&sC&\Lambda_{3}\end{matrix}\right]$
@@
where $M=\left[\begin{matrix}\Lambda_{2}+sQ&-sC^{T}\\ sC&\Lambda_{3}\end{matrix}\right]$ and $K=\left[\begin{matrix}Q&-C^{T}\\ C&0\end{matrix}\right].$ Since, ${\mathscr{P}}^{-1}_{LPESS}\mathcal{A}$ is similar to $\left[\begin{matrix}s^{-1}I&0\\ 0&M^{-1}K\end{matrix}\right],$ ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has eigenvalue $1/s$ with multiplicity $n$ and the remaining eigenvalues satisfies the generalized eigenvalue problem $Kp=\lambda Mp.$
@@
Let $A\in{\mathbb{R}}^{n\times n},\Lambda_{2}\in{\mathbb{R}}^{m\times m}$ and $\Lambda_{3}\in{\mathbb{R}}^{p\times p}$ be SPD matrices and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Assume that $s>0,$ then LPESS preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ has $n$ repeated eigenvalues equal to $1/s.$ The remaining $m+p$ eigenvalues satisfies the following:
@@ 1. $(1)$
    $\Big{[}\min\Big{\{}\frac{\vartheta_{\min}}{1+s\vartheta_{\min}},\frac{\tilde{\theta}_{\min}}{\vartheta_{\max}+s\tilde{\theta}_{\min}}\Big{\}},\frac{\vartheta_{\max}}{1+s\vartheta_{\max}}\Big{]}.$
@@ 2. $(2)$
    $\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\leq|\lambda|\leq\sqrt{\frac{\tilde{\theta}_{\max}}{1+s\vartheta_{\min}+s^{2}\tilde{\theta}_{\max}}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \frac{1}{s(1+s\sqrt{\tilde{\theta}_{\max}})}\leq|\lambda-\frac{1}{s}|\leq\frac{2}{s(2+s\vartheta_{\min})},$
where $\vartheta_{\max}:=\lambda_{\max}(\Lambda_{2}^{-1}Q),$ $\vartheta_{\min}:=\lambda_{\min}(\Lambda_{2}^{-1}Q)$ $\tilde{\theta}_{\max}:=\lambda_{\max}(\Lambda_{3}^{-1}C\Lambda_{1}^{-2}C^{T})$ and $\tilde{\theta}_{\min}:=\lambda_{\min}(\Lambda_{3}^{-1}C\Lambda_{1}^{-2}C^{T}).$
@@
where $\mathcal{F}=\left[\begin{matrix}I&0\\ 0&-I\end{matrix}\right],$ $\widetilde{M}=\left[\begin{matrix}\Lambda_{2}+sQ&sC^{T}\\ -sC&\Lambda_{3}\end{matrix}\right]$ and $\widetilde{K}=\left[\begin{matrix}Q&C^{T}\\ -C&0\end{matrix}\right].$ Hence, $M^{-1}K$ is similar to $\widetilde{M}^{-1}\widetilde{K}.$ Now, $\widetilde{M}^{-1}\widetilde{K}$ is in the form of the preconditioned matrix discussed in [42]. Therefore, by applying the Corollary 4.1 and Theorem 4.3 of [42], we obtain the desired bounds in $(1)$ and $(2).$blacksquare
6 The strategy of parameter selection
It is worth noting that the efficiency of the proposed PESS preconditioner depends on the selection of involved SPD matrices $\Lambda_{1},\Lambda_{2},\Lambda_{3}$ and a positive real parameter $s.$ In this section, motivated by [42, 43, 44], we discuss a practical way to choose the parameter $s$ and the SPD matrices for the effectiveness of the proposed preconditioners.
It is well acknowledged that the optimal parameter of PESS iteration process is obtained when $\vartheta(\mathcal{T})$ is minimized [43]. For achieving this, similar to the approach in [44], we first define a function $\varphi(s)=\|\mathcal{Q}_{PESS}\|^{2}_{F}$ depending on the parameter $s$. Our aim is to minimize $\varphi(s).$ Then after some straightforward calculations, we obtain
$\varphi(s)=\|\mathcal{Q}_{PESS}\|^{2}_{F}$ $=tr(\mathcal{Q}^{T}_{PESS}\mathcal{Q}_{PESS})$
$=\|\Lambda_{1}\|^{2}_{F}+\|\Lambda_{2}\|^{2}_{F}+\|\Lambda_{3}\|^{2}_{F}+(s-1)^{2}\|A\|^{2}_{F}+2(s-1)tr(\Lambda_{1}A)$
$\quad+2(s-1)^{2}\|B\|_{F}^{2})+2(s-1)^{2}\|C\|_{F}^{2}),$
where $tr(\cdot)$ and $\|\cdot\|_{F}$ denote the trace and the Frobenious norm of a matrix, respectively. Now we choose the parameter $s$ and SPD matrices $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ to make $\varphi(s)$ as small as possible. Since $\|A\|^{2}_{F},\|B\|^{2}_{F},\|C\|^{2}_{F}$ and $tr(\Lambda_{1}A)$ are positive, we can select $s=1,$ then $\varphi(s)=\|\Lambda_{1}\|^{2}_{F}+\|\Lambda_{2}\|^{2}_{F}+\|\Lambda_{3}\|^{2}_{F}.$ Thus if we choose $\|\Lambda_{1}\|_{2},\|\Lambda_{2}\|_{2},\|\Lambda_{3}\|_{2}\,\rightarrow 0,$ we have $\varphi(s)\rightarrow 0$ and consequently, $\mathcal{Q}_{PESS}\rightarrow 0.$
On the other hand, motivated by [45, 46], we discuss another strategy for choosing the parameter $s$. Notice that in Algorithm 1, we need to solve two linear system with coefficient matrices $\widehat{X}=\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C$ and $\widetilde{A}=\Lambda_{1}+sA+s^{2}B^{T}\widehat{X}B.$ Similar to [45, 46], we choose $s$ and $\|\Lambda_{2}\|_{2}$ as follows:
$s=\sqrt{\frac{\|\Lambda_{2}\|_{2}}{\|C^{T}\Lambda_{3}^{-1}C\|_{2}}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \|\Lambda_{2}\|_{2}=\frac{\|B\|_{2}^{4}}{{4\|C^{T}\Lambda_{3}^{-1}C\|_{2}\|A\|_{2}^{2}}},$ (6.1)
which balance the matrices $\Lambda_{2}$ and $C^{T}\Lambda_{3}^{-1}C$ in $\widehat{X}$ and the matrices $A$ and $C^{T}\Lambda_{3}^{-1}C$ in $\widetilde{A}$. In this case, we denote $s$ by $s_{est}$ and $\|\Lambda_{2}\|_{2}$ by $\beta_{est}.$ Numerical results are presented in Section 7 to demonstrate the effectiveness of ${\mathscr{P}}_{PESS}$ for the above choices of the parameters.
In the sequel, we can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where
$\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$
Since the prefactor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent. A general criterion for a preconditioner to perform efficiently is that it should closely approximate the coefficient matrix $\mathcal{A}$ [10]. Consequently, the difference $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}}$ approaches zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, the preconditioner PESS shows enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\widetilde{A}$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned. Similar investigations also hold for LPESS preconditioner. Nevertheless, in Figure 9, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$